# Real-time Neural Rendering of LiDAR Point Clouds 2Research Foundation - Flanders

### v1

@@ 2 Related Work
Neural Point Cloud Rendering The majority of methods train neural feature vectors on a per scene basis [hu2023trivolpointcloudrendering], or require encoding them using a pretrained network [Rakhimov_2022_CVPR]. Rendering involves projecting the feature-enriched points and reconstructing them using CNNs, typically a U-net. While some methods obtain features from the point cloud itself [hu2023point2pixphotorealisticpointcloud], others rely on a set of registered photographs from the same scene [npbg, franke2024trips, Rakhimov_2022_CVPR]. Because each point requires such features, these methods do not scale well with input size, which can easily run into dozens of millions of points for LiDAR scans. Our method foregoes any form of scene-specific training or encoding by directly rasterizing colored points into the output view, requiring only a U-net to reconstruct the final image.
Filtering Our depth-guided filter is designed to discard unwanted background information. A similar problem has been addressed in InvSfM [pittaluga2019revealing], where the authors train a dedicated U-net to resolve foreground visibility, but this comes at the cost of increased inference time and requires a suitable ground truth dataset. NPBG++ [Rakhimov_2022_CVPR] employs a depth heuristic similar to ours to transfer input image features to aggregated features on the point cloud. Theirs operates on only one level that has to set based on the chosen dataset, whereas ours is multi-resolution.
Direct Rendering Pointersect [pointersect] renders colored point clouds based on a model that is trained to predict ray intersection tests. This approach is scene agnostic and works on different types of point clouds, such as RGBD and LiDAR. The main drawback of this technique is its inference time, as rendering one image can take up to several minutes. It also does not address the difficulty of mitigating defects and gaps seen in LiDAR scans.
@@ 3 Method
U-Net Reconstruction The original U-Net [unet] focused on segmentation, where the goal was to segment biomedical images. Other works have shown that this architecture performs well for other image-processing tasks, such as denoising [gurrola2021residual]. We modify the U-Net to take RGBDA as input, where D represents z-buffer depth, and A (alpha) indicates whether a point was projected on this pixel. Points that were discarded by our depth-guided filtering will also have a zero alpha value. Instead of using cross-entropy loss combined with pixel-wise soft-max, we take inspiration from novel view synthesis [kerbl3Dgaussians] and combine loss and a perceptual similarity loss to balance pixel-level accuracy with perceptual similarity.
where and . We also added Batch normalization after each convolutional step during downsampling to accelerate training. We trained our network for 170 epochs on 36 ScanNet++ v1 scenes [yeshwanthliu2023scannetpp], consisting of 19,188 training image pairs and 4,798 test image pairs, where each pair consists of a ground-truth image and a rendered LiDAR point cloud image. We used random crops of pixels to reduce memory usage during training.
@@ 4 Experiments
Our method is implemented in CUDA, allowing for real-time rendering of point clouds. All experiments were run on an NVIDIA GeForce RTX 4090 and 13th Gen Intel(R) Core(TM) i9-13900KF. Training took around 60 hours on the same machine. We evaluated our method on a random subset of 55 scenes from ScanNet++ v2 and removed them from our training set. For comparison, we evaluate our method against Pointersect [pointersect]. The point clouds have an average size of 23 million points without any decimation. Results are shown in Table 1 and Fig. 3. We note that metrics in Table 1 are subject to slightly inaccurate poses of ground truth images in ScanNet++, and the color distribution of the ground truth images the point colors of the LiDAR scanner do not align because they originate from different cameras. Still, we present them here due because we believe they provide a reasonable indication of quality. We also compared reconstruction with and without the depth filter and noticed little to no difference in quality metrics. However, we notice a clear difference upon visual inspection. Using the depth filter tends to produce sharper details. It also includes more high-frequency, noisy artifacts than without the depth-guided filter. The latter does better at filling holes but tends to produce a more blurred and washed-out appearance. We find that using the depth filter results in clearer, more apprehensible images, especially for scenes with high depth complexity where background leaking is more prominent. A visual comparison can be found in the supplementary material.
@@ 5 Discussion & Conclusion
[…]

### v2

@@ 2 Related Work
Neural Point Cloud Rendering The majority of methods train neural feature vectors on a per scene basis [1], or require encoding them using a pretrained network [2]. Rendering involves projecting the feature-enriched points and reconstructing them using CNNs, typically a U-net. While some methods obtain features from the point cloud itself [3], others rely on a set of registered photographs from the same scene [4, 5, 2]. Because each point requires such features, these methods do not scale well with input size, which can easily run into dozens of millions of points for LiDAR scans. Our method foregoes any form of scene-specific training or encoding by directly rasterizing colored points into the output view, requiring only a U-net to reconstruct the final image.
Filtering Our depth-guided filter is designed to discard unwanted background information. A similar problem has been addressed in InvSfM [6], where the authors train a dedicated U-net to resolve foreground visibility, but this comes at the cost of increased inference time and requires a suitable ground truth dataset. NPBG++ [2] employs a depth heuristic similar to ours to transfer input image features to aggregated features on the point cloud. Theirs operates on only one level that has to set based on the chosen dataset, whereas ours is multi-resolution.
Direct Rendering Pointersect [7] renders colored point clouds based on a model that is trained to predict ray intersection tests. This approach is scene agnostic and works on different types of point clouds, such as RGBD and LiDAR. The main drawback of this technique is its inference time, as rendering one image can take up to several minutes. It also does not address the difficulty of mitigating defects and gaps seen in LiDAR scans.
@@ 3 Method
U-Net Reconstruction The original U-Net [8] focused on segmentation, where the goal was to segment biomedical images. Other works have shown that this architecture performs well for other image-processing tasks, such as denoising [9]. We modify the U-Net to take RGBDA as input, where D represents z-buffer depth, and A (alpha) indicates whether a point was projected on this pixel. Points that were discarded by our depth-guided filtering will also have a zero alpha value. Instead of using cross-entropy loss combined with pixel-wise soft-max, we take inspiration from novel view synthesis [10] and combine loss and a perceptual similarity loss to balance pixel-level accuracy with perceptual similarity.
where and . We also added Batch normalization after each convolutional step during downsampling to accelerate training. We trained our network for 170 epochs on 36 ScanNet++ v1 scenes [11], consisting of 19,188 training image pairs and 4,798 test image pairs, where each pair consists of a ground-truth image and a rendered LiDAR point cloud image. We used random crops of pixels to reduce memory usage during training.
@@ 4 Experiments
Our method is implemented in CUDA, allowing for real-time rendering of point clouds. All experiments were run on an NVIDIA GeForce RTX 4090 and 13th Gen Intel(R) Core(TM) i9-13900KF. Training took around 60 hours on the same machine. We evaluated our method on a random subset of 55 scenes from ScanNet++ v2 and removed them from our training set. For comparison, we evaluate our method against Pointersect [7]. The point clouds have an average size of 23 million points without any decimation. Results are shown in Table 1 and Fig. 3. We note that metrics in Table 1 are subject to slightly inaccurate poses of ground truth images in ScanNet++, and the color distribution of the ground truth images the point colors of the LiDAR scanner do not align because they originate from different cameras. Still, we present them here due because we believe they provide a reasonable indication of quality. We also compared reconstruction with and without the depth filter and noticed little to no difference in quality metrics. However, we notice a clear difference upon visual inspection. Using the depth filter tends to produce sharper details. It also includes more high-frequency, noisy artifacts than without the depth-guided filter. The latter does better at filling holes but tends to produce a more blurred and washed-out appearance. We find that using the depth filter results in clearer, more apprehensible images, especially for scenes with high depth complexity where background leaking is more prominent. A visual comparison can be found in the supplementary material.
@@ 5 Discussion & Conclusion
Acknowledgements This research was partly funded by the Special Research Fund (BOF) of Hasselt University (R-14360, R-14436) and the FWO fellowship grant (1SHDZ24N), the European Union (HORIZON MAX-R, Mixed Augmented and Extended Reality Media Pipeline, 101070072), the Flanders Make’s XRTwin SBO project (R-12528). This work was made possible with support from MAXVR-INFRA, a scalable and flexible infrastructure that facilitates the transition to digital-physical work environments.